OpenStudy (anonymous):

evaluate the improper integral if it is convergent: ∫1/((2x-1)^(3/2)) dx from 0 to ∞

OpenStudy (anonymous):

from 1 to ∞, sorry

OpenStudy (anonymous):

\[\int_1^{\infty } \frac{1}{(2 x-1)^{3/2}} \, dx\]

OpenStudy (anonymous):

quick question... how did u make the integral sign with the limits?

OpenStudy (anonymous):

\int_1^{\infty } \frac{1}{(2 x-1)^{3/2}} \, dx

OpenStudy (anonymous):

put between \[

OpenStudy (anonymous):

\[\int_1^r \frac{1}{(2 x-1)^{3/2}} \, dx\]

OpenStudy (anonymous):

i got u=2x-1 and du=2dx

OpenStudy (anonymous):

using u sub , it evaluate to \[1-\frac{1}{\sqrt{-1+2 r}}\]

OpenStudy (anonymous):

as r goes to infinity, second term goes to 0. Leaving only 1

OpenStudy (anonymous):

can you use b instead of r please?

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

how did you get 1-1/(√-1+2b)?

OpenStudy (anonymous):

u=2x-1 du= 2 dx

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

i got the lim b -> ∞ is -1/(√2b-1) + 1 which I think is what you said

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

\[\frac{1}{2} \int \frac{1}{u^{3/2}} \, du\] \[-\frac{1}{\sqrt{u}}\]

OpenStudy (anonymous):

yep got that

OpenStudy (anonymous):

\[-\frac{1}{\sqrt{2x-1}}\]

OpenStudy (anonymous):

does x = b?

OpenStudy (anonymous):

i got what you got but -1 at the end

OpenStudy (anonymous):

because i had to subtract when u=1 from when u=2b-1

OpenStudy (anonymous):

\[-\frac{1}{\sqrt{2b-1}}-\left(-\frac{1}{\sqrt{2(1)-1}}\right)\]

OpenStudy (anonymous):

\[1-\frac{1}{\sqrt{-1+2 b}}\]

OpenStudy (anonymous):

\[\lim_{b\to \infty } \, \left(1-\frac{1}{\sqrt{2 b-1}}\right)\]

OpenStudy (anonymous):

okay im following

OpenStudy (anonymous):

as b approaches infinity , the second term goes to 0

OpenStudy (anonymous):

the second term goes to 0 right? because the number you are dividing x by is getting greater and greater

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!!!